# SemiHMER: Semi-supervised Handwritten Mathematical Expression Recognition using pseudo-labels

### v1

@@ Abstract
In recent years, deep learning with Convolutional Neural Networks (CNNs) has achieved remarkable results in Handwritten Mathematical Expression Recognition (HMER). However, improving the performance with limited labeled training data remains challenging. This paper presents, for the first time, a simple yet effective semi-supervised HMER framework by introducing dual-branch semi-supervised learning. Specifically, we simplify the conventional deep co-training from consistency regularization to cross-supervised learning, in which the prediction of one branch is used as the pseudo-labels to supervise the other branch directly end-to-end. Considering that the learning of the two branches tends to converge in the later stages of model optimization, we also incorporate a weak-to-strong strategy by applying different levels of augmentation to each branch, which expands the training data and improves the training quality of the network. Meanwhile, to enhance the performance of the HMER decoder, a novel module called Global Dynamic Counting Module(GDCM) is proposed, which could alleviate inaccuracies of recognition in long-distance formula recognition and the occurrence of repeated characters. The code of our model will be released at https://github.com/chenkehua/SemiHMER.
@@
Handwritten Mathematical Expression Recognition, Semi-supervised Learning, Weak-to-Strong Augmentation, Global Dynamic Counting Module
@@ 1 Introduction
Handwritten mathematical expression recognition (HMER) is a crucial task in the field of Optical Character Recognition (OCR), and it has been widely applied in digital libraries, office automation, automatic grading, and question parsing. In recent years, HMER has achieved significant success with the advancement of artificial intelligence. However, due to the complex structure of formulas, personalized handwriting styles, and the substantial cost and time involved in manual labeling, HMER remains a challenging task.
@@
Consistency regularization-based methods have been widely used in semi-supervised learning to effectively utilize unlabeled images, which promotes the network to generate similar predictions for the same unlabeled images with different augmentations by calculating the differences between outputs as a loss function. They typically include three methods: input perturbation, feature perturbation, and network perturbation of input images.
@@
Self-training has also been employed in semi-supervised learning research. It starts by training an initial model on labeled images to obtain pseudo-labels for unlabeled images, incorporating them to expand the training data, and then retraining the model to get boosted. One approach is called cross pseudo-supervision, which feeds both labeled and unlabeled images into two networks that share the same structure but have different parameter initializations, adding constraints to ensure similar outputs from both networks for the same sample. Here, we propose a more novel and simpler consistency regularization method where we apply pseudo-supervision on unlabeled images, which serves as an additional signal to supervise another branch network. Motivated by FixMatch’s approach[18] of imposing consistency constraints between predictions generated from weak and strong augmentations, we develop a weak-to-strong cross-head collaborative training framework for HMER tasks, achieving higher accuracy than conventional supervised training structures that rely solely on labeled data. This approach effectively expands the training data, leading to enhanced training quality of the recognition network.
@@
Meanwhile, we propose GDCM(Global Dynamic Counting Module) to alleviate the misrecognition issues in long-distance formulas and the misrecognition of repeated identical characters. While updating the hidden state and context information at each step, GDCM dynamically updates the obtained global counting vector based on the prediction from the previous time step. The updated counting vector is then used to compute the output probability at the current time step.
The main contributions of this paper can be summarized as follows:
@@
    We introduce an innovative weak-to-strong augmentation strategy designed to enhance the performance of pseudo-supervision learning. By progressively applying stronger image augmentation techniques, our approach not only improves the robustness of the model but also maximizes the utilization of training data, which leads to significant advancements in recognition accuracy.
@@ 2.2 Semi-supervised learning
Semi-supervised object detection. STAC[19] proposed a semi-supervised object detection algorithm based on hard pseudo labels. Ambiguity-Resistant Semi-supervised Learning (ARSL)[14], including Joint-Confidence Estimation (JCE) and Task-Separation Assignment (TSA), is generally applicable to single-stage semi-supervised object detection tasks. ISTM[32] proposed an interactive self-training model. It integrates object detection results from different iterations using Non-Maximum Suppression (NMS) and uses two ROI heads with different structures to estimate each other’s pseudo-labels. TLDR[23] proposed a new semi-supervised object detector called Consistent-Teacher by analyzing the existing issues of pseudo-label misalignment and inconsistency in semi-supervised object detection.
@@ 3.1 Overview
Fig.1 shows the overview of our cross-decoder architecture, in which labeled and unlabeled images are fed into a shared CNN network. Following CAN[11], we apply the DenseNet as the backbone to extract features. Instead of using separate individual models, the model employs two identical decoder heads called cross-head, which enables the learning of more compact features and further enhances the model’s generalization capability.
Between ground truth and the corresponding predicted pseudo-labels, as for labeled images, the pseudo-label is generated by the prediction from a weakly augmented image. Then, this pseudo-label is used to supervise the prediction of a strongly augmented image from another head. Similarly, as for unlabeled images, we only use the pseudo-labels generated from the weakly augmented branch to supervise the prediction results of the strongly augmented branch. On the one hand, the pseudo-label plays a significant role in expanding the training data. On the other hand, imposing constraints on the predictions of weak and strong images can benefit from the advantages of consistency training. In addition, as shown in Fig.1, week-to-strong augmentation and dynamic branch training strategy boost consistency training samples’ diversity while dealing with longer texts and the long-tail data of complex characters.
The proposed approach is composed of a shared backbone $F$ and two parallel decoders $G_{m}(m\in\{1,2\})$, with both decoders having identical structures. The pseudo label is generated by $P_{m}(y|x)=G_{m}(F(x))$, which is produced by $F$ and $Gm$. The pseudo label will then be used as the supervision signal for the other decoder.
Figure 1: Overview of our Semi-supervised Handwritten Mathematical Expression Recognition using pseudo-labels. In this scheme, we use ${Decoder_{1}}$ as an example. Weakly augmented (or even non-augmented) labeled data flows through the encoder module and the corresponding prediction ${Decoder_{1}}$ to produce the prediction ${pred_{1}}$, which is then supervised by the ground truth. At the same time, the prediction ${pred_{1}}$ is also utilized as a supervision signal for the strongly augmented image prediction ${pred_{2}}$ from the other head, and vice versa. Besides that, We use predictions from weakly augmented unlabeled data as ground truth to supervise the encoder and predictions on strongly augmented data
@@ 3.2 Weak-to-strong augmentation
To fully leverage the advantages of consistency training, We utilize weak and strong augmentations to introduce additional information into our framework. In our experiments, weak augmentation means little to no augmentation, whereas strong augmentation combines distortions, stretching, and perspective changes. Specifically, we randomly distort images every time, stretch images with a probability of 50%, and perspective images with a probability of 30%.
@@ 3.3 Global Dynamic Counting Module
To overcome these limitations and better leverage the relationship between formulas and count vectors to guide the model, as shown in Fig.2, we introduce the concept of GDCM(Global Dynamic Counting Module) into the Handwritten Mathematical Expression Recognition task. Specifically, based on the encoder-decoder recognition logic, the decoder relies on the previous state to predict the current output at each time step. Therefore, the feature information input to the decoder should be dynamically updated at each step. Our model achieves this by dynamically updating the global count vector based on the predicted value from the previous time step, while also updating the hidden state and context information. The updated count vector is then used to calculate the output probability for the current time step.
@@ 3.4 Loss function
In the setting of semi-supervised learning, we provide a batch of labeled samples $\mathcal{D}_{l}$ at each iteration. We preliminarily $\mathcal{L}_{cls}$ as standard cross entropy classification loss of the predicted probability $\mathcal{P}_{m}(y_{t})$, with $\mathcal{W}$ and $\mathcal{S}$ representing the corresponding weak and strong augmentation functions applied to the image respectively. The $\mathcal{L}_{cls}^{unsup}$ defined as:
$\mathcal{L}_{cls}^{unsup}=\mathcal{L}_{cls}^{\mathcal{W}}\mathcal{P}(y_{t})+\mathcal{L}_{cls}^{\mathcal{S}}\mathcal{P}(y_{t})$ (4)
where $\mathcal{L}_{sup}$ was defined as :
@@
For the unlabeled samples $\mathcal{D}_{un}$ provided at each iteration, we also use the standard cross-entropy classification loss for supervision, where the predictions from weak supervision are used to supervise the predictions from strong supervision, defined as $\mathcal{L}_{cls}^{unsup}$ as follows:
@@
The counting ground truth for each symbol class is represented by $\widehat{\mathcal{V}}$ is a smooth L1 loss[17] function defined as follows:
$\begin{split}\mathcal{L}_{counting}=&Smooth\mathcal{L}1\cdot\mathcal{W}(\mathcal{V},\widehat{\mathcal{V}})\\ &+Smooth\mathcal{L}1\cdot\mathcal{S}(\mathcal{V},\widehat{\mathcal{V}})\end{split}$ (7)
The total loss function is composed of two components and is defined in the following manner:
@@ 4.1 Datasets
Figure 3: Some example images from the CROHME dataset
@@
CROHME Dataset[16] is from the competition on recognition of online handwritten mathematical expression, which is the most widely used public dataset. It is derived from the competition focused on recognizing online handwritten mathematical expressions. The CROHME dataset creates images from handwritten stroke trajectory data found in InkML files, leading to images with clean backgrounds. The CROHME training set number is 8,836, while the test set contains 986, 1147 and 1199 images, including 111 symbol classes (C), including ”sos” and ”eos”. Some example images are shown in Fig.3. We use 2,000 images synthesized in CROHME2023[29] as unlabeled images that exhibit similar handwriting styles and share a consistent character set.
@@ 4.2 Implementation details
The proposed SemiHMER is implemented in PyTorch. To accelerate the iteration efficiency, we used four Nvidia Tesla V100 GPUs, each equipped with 32GB RAM, to train the model. The batch size is set to 12, and the Adadelta[34] optimizer is used for training. The learning rate starts at 0 and monotonically increases to 2 by the end of the first epoch. After that, the learning rate decays to 0 according to a cosine annealing schedule. To enhance the effectiveness of semi-supervised learning, we did not use cross-training in the first 240 epochs, allowing each branch to train independently. After 240 epochs, we conducted semi-supervised cross-training with a learning rate of 0.2 and further fine-tuned for another 240 epochs. Our method relies on data augmentation techniques, and we employed various augmentation methods (distortion, affine transformation, and stretching) to explore the capabilities of our approach.
@@ 4.3 Evaluation Metrics
Table 1: Expression Recognition Rate (ExpRate) performance of SemiHMER and other state-of-the-art methods on CROHME 2014, CROHME 2016 and CROHME 2019 test set
@@
CAN-aug 58.32 75.05 81.54 57.28 74.11 81.34 57.63 75.15 82.24
SemiHMER 60.95 76.27 82.25 58.23 73.93 81.51 58.63 76.98 83.15
4.4 Comparison with SOTA
To demonstrate the superiority of SemiHMER, we compare it with previous state-of-the-art (SOTA) methods. Since SemiHMER heavily relies on the effectiveness of data augmentation, we primarily focus on the results obtained when data augmentation is included during the model training process. As shown in Table 1, using CAN as the baseline, our multi-GPU training with four V100 GPUs and a learning rate of 2 did not achieve the original performance of CAN, yielding only 55.48(CROHME2014) which is 1.52% lower than the 57.00 proposed by CAN. After applying data augmentation, the performance improved by 2.92%. With the aid of optimal data augmentation, dual-branch semi-supervision learning further enhanced the model by 2.55%. This indicates that our method can effectively leverage data augmentation to significantly boost the model’s performance.
@@ 4.5 Ablation Study
4.5.1 Effectiveness of GDCM
Table 2: Alabtion study of GDCM
GDCM Results in CROHME dataset
2014 2016 2019
55.48 53.36 53.38
56.30 56.14 53.54
As shown in Table 2, GDCM boosts performance to 56.30%, 56.14%, and 53.54% by updating the hidden state and context information, which shows the effectiveness of our Global Dynamic Counting method.
4.5.2 Strategy of the Weak-to-Strong augmentation
By utilizing the weak-strong strategy, we can introduce more information to improve consistency. The strong transformation function generates augmented images with different intensities each time, which leads to improved model performance. However, excessive augmentation can cause a performance decline, possibly due to a significant increase in erroneous predictions from the other branch, misleading the network’s optimization direction.
Table 3: Ablation study of different augmentations strategy for pseudo supervision learning strategy
Strategy Results in CROHME dataset
distort stretch perspective semi-supervision CROHME2014 CROHME2016 CROHME2019

### v3

@@ Abstract
In this paper, we study semi-supervised Handwritten Mathematical Expression Recognition (HMER) via exploring both labeled data and extra unlabeled data. We propose a novel consistency regularization framework, termed SemiHMER, which introduces dual-branch semi-supervised learning. Specifically, we enforce consistency between the two networks for the same input image. The pseudo-label, generated by one perturbed recognition network, is utilized to supervise the other network using the standard cross-entropy loss. The SemiHMER consistency encourages high similarity between the predictions of the two perturbed networks for the same input image and expands the training data by leveraging unlabeled data with pseudo-labels. We further introduce a weak-to-strong strategy by applying different levels of augmentation to each branch, effectively expanding the training data and enhancing the quality of network training. Additionally, we propose a novel module, the Global Dynamic Counting Module (GDCM), to enhance the performance of the HMER decoder by alleviating recognition inaccuracies in long-distance formula recognition and reducing the occurrence of repeated characters. The experimental results demonstrate that our work achieves significant performance improvements, with an average accuracy increase of 5.47% on CROHME14, 4.87% on CROHME16, and 5.25% on CROHME19, compared to our baselines.
@@
Handwritten Mathematical Expression Recognition Semi-supervised Learning Global Dynamic Counting Module.
@@ 1 Introduction
Handwritten mathematical expression recognition (HMER) is a crucial task in Optical Character Recognition (OCR) with applications in digital libraries, office automation, automatic grading. Despite significant progress, HMER remains challenging due to the complex structure of formulas, personalized handwriting styles, and the substantial cost and time involved in manual labeling. Semi-supervised learning, which leverages both labeled and unlabeled data, offers a promising solution to these challenges.
@@
In semi-supervised learning, consistency regularization and self-training are two widely used techniques[18]. Consistency regularization-based methods promote the network to generate consistent predictions for the same input with different augmentations by calculating the differences between outputs as a loss function. This technique effectively utilizes unlabeled data through three main approaches: input perturbation (e.g., applying transformations like rotation or scaling to the input image), feature perturbation (e.g., introducing noise or variations in the feature space), and network perturbation (e.g., using different network architectures or parameter initializations to generate diverse predictions).
@@
Self-training has also been employed in semi-supervised learning research. It starts by training an initial model on labeled images to obtain pseudo-labels for unlabeled images, incorporating them to expand the training data, and then retraining the model to get boosted. One approach is called cross pseudo-supervision, which feeds both labeled and unlabeled images into two networks that share the same structure but have different parameter initializations, adding constraints to ensure similar outputs from both networks for the same sample. Here, we propose a more novel and simpler consistency regularization method where we apply pseudo-supervision on unlabeled images, which serves as an additional signal to supervise another branch network.
Motivated by FixMatch’s approach[20] of imposing consistency constraints between predictions generated from weak and strong augmentations, we develop a weak-to-strong cross-head collaborative training framework for HMER tasks, achieving higher accuracy than conventional supervised training structures that rely solely on labeled data. In our framework, weak augmentation refers to applying minimal even no image transformation, while strong augmentation involves applying diverse and extensive transformations (e.g., distortion, stretch, perspective). Our weak-to-strong augmentation strategy applies weak augmentation to one branch and strong augmentation to the other, intentionally creating a significant divergence in their predictions. By enforcing consistency regularization between these two branches, we ensure that the model learns robust representations from the same input image under different augmentation conditions. This approach effectively expands the training data, leading to enhanced training quality of the recognition network.
@@
Meanwhile, we propose a GDCM (Global Dynamic Counting Module) to alleviate the misrecognition issues in long-distance formulas and the misrecognition of repeated identical characters. While updating the hidden state and context information at each step, GDCM dynamically updates the obtained global counting vector based on the prediction from the previous time step. The updated counting vector is then used to compute the output probability at the current time step. The main contributions of this paper can be summarized as follows.
@@
    We introduce an innovative weak-to-strong augmentation strategy designed to enhance the performance of pseudo-supervision learning. By progressively applying stronger image augmentation techniques, our approach not only improves model robustness but also maximizes the utilization of training data, leading to significant advancements in recognition accuracy.
@@ 2.2 Semi-supervised learning
Semi-supervised object detection. STAC[21] proposed a semi-supervised object detection algorithm based on hard pseudo-labels. Ambiguity-Resistant Semi-supervised Learning (ARSL)[15], including Joint-Confidence Estimation (JCE) and Task-Separation Assignment (TSA), is generally applicable to single-stage semi-supervised object detection tasks. ISTM[34] proposed an interactive self-training model. It integrates object detection results from different iterations using Non-Maximum Suppression (NMS) and uses two ROI heads with different structures to estimate each other’s pseudo-labels. TLDR[25] proposed a new semi-supervised object detector called Consistent-Teacher by analyzing the existing issues of pseudo-label misalignment and inconsistency in semi-supervised object detection.
@@ 3.1 Overview
Figure 1: Overview of our Semi-supervised Handwritten Mathematical Expression Recognition using pseudo-labels. In this scheme, we use ${Decoder_{1}}$ as an example. Weakly augmented (non-augmented) labeled data flows through the encoder module and the corresponding prediction ${Decoder_{1}}$ to produce the prediction ${Classification\ Vector_{super,weak}}$, which is then supervised by the ground truth. At the same time, the prediction ${Classification\ Vector_{super,weak}}$ will be transformed into hard pseudo labels, which also utilized as a supervision signal for the strongly augmented image prediction ${Classification\ Vector_{super,strong}}$ from the other head. The ${Classification\ Vector_{super,strong}}$ is also supervised by the ground truth labels. Besides that, We use predictions ${Classification\ Vector_{unsuper,weak}}$ from weakly augmented unlabeled data as ground truth to supervise the encoder and predictions ${Classification\ Vector_{unsuper,strong}}$ on strongly augmented data. Augmentation alternation refers to alternately switching the intensity of augmentation applied to different decoders.
Figure 1 shows our proposed semi-supervised framework for handwritten mathematical expression recognition. The framework employs a cross-decoder architecture that processes both labeled and unlabeled data through a shared network. Following CAN[12], we adopt DenseNet[9] as the backbone for feature extraction. Instead of using separate individual models, we innovatively design two identical decoder heads working in parallel, which we term the cross-head structure. This design enables more efficient feature learning and enhances the model’s generalization capability through mutual supervision.
Our semi-supervised learning strategy operates in two parallel streams. For labeled data processing, we first feed weakly augmented images through the encoder and one of the decoders to generate initial predictions. These predictions serve dual purposes: they are supervised by ground truth labels for word prediction and transformed into hard pseudo-labels to guide the other decoder’s predictions on strongly augmented versions of the same images. This cross-supervision mechanism ensures consistency between different augmentation views while maintaining alignment with ground truth. For unlabeled data, we employ a similar but simplified approach: the predictions from weakly augmented images serve as pseudo-labels to supervise the predictions of their strongly augmented counterparts, effectively leveraging unlabeled data through consistency training.
The proposed approach is composed of a shared backbone $F$ and two parallel decoders $G_{m}(m\in\{1,2\})$, with both decoders having identical structures. The pseudo label is generated by $P_{m}(y|x)=G_{m}(F(x))$, where $F$ is the backbone network and $Gm$ is the respective decoder. This pseudo-label is subsequently used as the supervision signal for the other decoder, creating a feedback loop that enhances learning.
@@ 3.2 Weak-to-strong augmentation
To fully leverage the advantages of consistency training, We utilize weak and strong augmentations to introduce additional information into our framework. In our experiments, weak augmentation means no augmentation, whereas strong augmentation combines distortions, stretching, and perspective changes. Specifically, we randomly distort images every time, stretch images with a probability of 50%, and perspective images with a probability of 30%.
@@
By alternating the application of augmentations and dynamically training the decoders, our framework not only maximizes the effective use of both labeled and unlabeled data but also fosters a robust learning environment that is capable of accurately recognizing handwritten mathematical expressions. This systematic approach ultimately leads to improved performance in real-world applications.
@@ 3.3 Global Dynamic Counting Module
To overcome these limitations and better leverage the relationship between formulas and count vectors to guide the model, we introduce the concept of GDCM(Global Dynamic Counting Module) into the Handwritten Mathematical Expression Recognition task. Specifically, based on the encoder-decoder recognition logic, the decoder relies on the previous state to predict the current output at each time step. Therefore, the feature information input to the decoder should be dynamically updated at each step. Our model achieves this by dynamically updating the global count vector based on the predicted value from the previous time step, while also updating the hidden state and context information. The updated count vector is then used to calculate the output probability for the current time step.
@@ 3.4 Loss function
In the setting of semi-supervised learning, we provide a batch of labeled samples $\mathcal{D}_{l}$ at each iteration. We preliminarily $\mathcal{L}_{cls}$ as standard cross entropy classification loss of the word predicted probability $\mathcal{P}(y_{t})$, with $\mathcal{W}$ and $\mathcal{S}$ representing the corresponding weak and strong augmentation functions applied to the image respectively. The $\mathcal{L}_{cls}^{unsup}$ defined as:
$\mathcal{L}_{cls}^{unsup}=\mathcal{L}_{cls}(\mathcal{P}^{\mathcal{S}}(y_{t}),\mathcal{P}^{\mathcal{W}}(y_{t}))$ (4)
where $\mathcal{L}_{cls}^{sup}$ was defined as :
@@
For the unlabeled samples $\mathcal{D}_{un}$ provided at each iteration, we also use the standard cross-entropy classification loss for supervision. Specifically, we denoted the unlabeled samples’s predictions as $y_{t}^{{}^{\prime}}$. The predictions from weak supervision are used to supervise the predictions from strong supervision, defined as $\mathcal{L}_{cls}^{unsup}$ as follows:
@@
Furthermore, for each decoder, there is a counting ground truth loss for each symbol class of decoder’s prediction result, represented by $\widehat{\mathcal{V}}$. It is a smooth L1 loss[19] function used to calculate the loss between the predicted count and the ground truth for each formula (not shown in 1), defined as follows:
$\mathcal{L}_{counting}=Smooth\mathcal{L}_{1}\cdot\mathcal{W}(\widehat{\mathcal{V}}_{1})+Smooth\mathcal{L}_{1}\cdot\mathcal{S}(\widehat{\mathcal{V}}_{2})$ (7)
So the total loss function is composed of four components and is defined in the following manner:
@@ 4.1 Datasets
Figure 3: Some example images from the CROHME dataset.
@@
CROHME Dataset[17] is from the competition on recognition of online handwritten mathematical expression, which is the most widely used public dataset. It is derived from the competition focused on recognizing online handwritten mathematical expressions. The CROHME dataset creates images from handwritten stroke trajectory data found in InkML files, leading to images with clean backgrounds. The CROHME training set number is 8,836, while the test set contains 986, 1147 and 1199 images, including 111 symbol classes (C), including "sos" and "eos". Some example images are shown in Fig. 2. We randomly selected 1,000 images from the CROHME 2023 dataset[31] as unlabeled data, with all characters included in the CROHME Dataset.
@@ 4.2 Implementation details
The proposed SemiHMER is implemented in PyTorch. To accelerate the iteration efficiency, we used four Nvidia Tesla V100 GPUs, each equipped with 32GB RAM, to train the model. It is worth noting that training SemiHMER on a single GPU would take at least one week, which is highly inefficient. The batch size is set to 12, and the Adadelta[36] optimizer is used for training. The learning rate starts at 0 and monotonically increases to 2 by the end of the first epoch. After that, the learning rate decays to 0 according to a cosine annealing schedule. To enhance the effectiveness of semi-supervised learning, we did not use cross-training in the first 240 epochs, allowing each branch to train independently. After 240 epochs, we conducted semi-supervised cross-training with a learning rate of 0.2 and further fine-tuned for another 240 epochs. Our method relies on data augmentation techniques, and we employed various augmentation methods (distortion, affine transformation, and stretching) to explore the capabilities of our approach.
@@ 4.3 Evaluation Metrics
Table 1: Expression Recognition Rate (ExpRate) performance of SemiHMER and other state-of-the-art methods on CROHME 2014, CROHME 2016, and CROHME 2019 test set.
@@
baseline-aug 58.32 75.05 81.54 57.28 74.11 81.34 57.63 75.15 82.24
SemiHMER 60.95(+5.47) 76.27 82.25 58.23(+4.87) 73.93 81.51 58.63(+5.25) 76.98 83.15
4.4 Comparison with State-of-the-Art
To demonstrate the superiority of SemiHMER, we compare it with previous state-of-the-art methods. Since SemiHMER heavily relies on the effectiveness of data augmentation, we primarily focus on the results obtained when data augmentation is included during the model training process. As shown in Table. 1, using CAN as the baseline, our multi-GPU training with four V100 GPUs and a learning rate of 2 did not achieve the original performance of CAN, yielding only 55.48% on CROHME 2014, 53.36% on CROHME 2016 and 53.38% on CROHME 2019 which are lower than the results proposed by CAN. However, despite extensive tuning of hyperparameters and configurations, we were unable to fully close this gap, highlighting a limitation of our current implementation. After applying data augmentation, the performance improved by 2.84%(58.32%) on CROHME 2014, 3.84%(57.2%) on CROHME 2016 and 4.25%(57.63%) on CROHME 2019. With the aid of optimal data augmentation, dual-branch semi-supervised learning further enhanced the model by 2.63%(60.95%) on CROHME 2014, 1.03%(58.23%) on CROHME 2016 and 1.0%(58.63%) on CROHME 2019 indicates that our method can effectively leverage data augmentation significantly boost the model’s performance.
@@ 4.5 Ablation Study
4.6 Effectiveness of Global Dynamic Counting Module
Table 2: Alabtion study of Global Dynamic Counting Module.
Method Result
CROHME 2014 CROHME 2016 CROHME 2019
no GDCM 55.48 53.36 53.38
with GDCM 56.30 56.14 53.54
GDCM boosts performance to 56.30,56.14,53.54 by updating the hidden state and context information, which shows the effectiveness of our Global Dynamic Counting method.
4.7 Effectiveness of the weak-to-strong augmentation strategy
By utilizing the weak-strong strategy, we can introduce more information to improve consistency. The strong transformation function generates augmented images with different intensities each time, which leads to improved model performance. However, excessive augmentation can cause a performance decline, possibly due to a significant increase in erroneous predictions from the other branch, misleading the network’s optimization direction. To allow the model to benefit from both data augmentation and cross-pseudo supervision strategies, we conducted experiments with different augmentation intensities. A natural idea is to gradually increase the intensity of augmentation. As shown in Table 3, by progressively introducing warping, stretching, and perspective transformations into the model, performance on the dataset CROHME2014 improved by 0.09%, 0.58%, and 2.02% respectively. Finally, when applying both weak-strong augmentations and cross-pseudo supervision strategies to both branches, the results further improved by 1.92%, 2.13%, and 2.63%, demonstrating the effectiveness of our strategy: as the augmentation becomes more effective, the impact of CPS becomes more pronounced.
Table 3: Ablation study of different augmentations strategy for pseudo supervision learning strategy.
Strategy Result
distort stretch perspective cps CROHME 2014 CROHME 2016 CROHME 2019